Displaying reports 861-880 of 1079.Go to page Start 40 41 42 43 44 45 46 47 48 End
Reports until 10:51, Monday 22 January 2024
GQuEST General
Daniel Grass - posted 10:51, Monday 22 January 2024 - last comment - 11:54, Tuesday 30 January 2024(11430)
Simple COMSOL Model of High Frequency Coating Thermal Noise

I using Levin's method (Internal thermal noise in the LIGO test masses: A direct approach) of using the suceptability of a mirror to a force resembling the laser beam profile to model the thermal noise, I modeled the GQuEST end mirrors as 2 mm thick, 1 in side length squares made of silicon with a Q of 10^6. I added a coating 20 um thick made of silicon as well but with a Q of 10^3. I constrained the 4 side of the mirror to only move in the z-axis. The spot radius is 2 mm. The face of the mirror had 52 x 52 mesh points, the substrate had 10 layers of mesh, and the coating had 3 layers of mesh. I simulated from 500 kHz to 16.5 MHz with 100 kHz frequency spacing. It look 5 hours and 17 minutes. Attached is the data converted into an ASD. The coating thermal noise, due to its low Q and that the force is applied on it, dominates the "bottom of the bucket".

Images attached to this report
Comments related to this report
Daniel Grass - 15:45, Thursday 25 January 2024 (11436)

I repeated the simulation but with a loss angle of 1e-4 instead of 1e-3 for the coating, no clamping of any kind, and a mesh 2x as dense in all 3 axes. The coating thermal noise does not fall off as hypothesized.

Images attached to this comment
Daniel Grass - 13:26, Friday 26 January 2024 (11437)

I repeated the most recent simulation but made the mesh 3x more coarse in all 3 axes. The amplitudes between the peaks is very similar. Interestingly, the frequencies of the peaks are changed compared to the fine mesh.

Images attached to this comment
Daniel Grass - 11:54, Tuesday 30 January 2024 (11443)

I added a back coating to the mirror. The loss angle for the coatings is 1e-4 and there is no clamping, but otherwise this simulation is identical to the original simulation.

Images attached to this comment
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 10:41, Monday 22 January 2024 (11429)
Newport Flexure Mount Noise Test in Michelson Interferometer

[Torrey, Daniel]

We set up a Michelson Interferometer using a Flexure Mirror Mount and a standard tip-tilt mirror mount as end mirrors to test if there was a resonance that produced the noise source around 520 Hz that is observed in the output filter cavity. The beamsplitter was a 50:50 non-polarizing plate, and the readout was done by a Thorlabs PDA50B2 amplified photodetector. The alignment was such that the contrast from hitting the table was above 80%. On the frequency trace, there was no discernable increase in noise around 520 Hz (see photos). I also played sinusoids at 520 Hz and did not observe any increase in the noise. This indicates this noise is not inherent to all flexure mounts. Perhaps some flexure mounts are manufactuered or mounted sufficiently differently as to produce a 520 Hz noise source. The large peaks appear to be 60 Hz noise and its overtones.

Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 14:40, Friday 19 January 2024 - last comment - 16:44, Friday 19 January 2024(11427)
mystery resonance investigation (cont)

[Jeff, Torrey]

This is a continuation from 11423. We set to improve upon our characterization of how well pushing a viton ring against the mirror (M2) of the cavity damped out the mystery 520 Hz resonance. 

We first drove sound from a phone at 490 Hz, 520 Hz, 550 Hz with and without pushing on the mirror, seen in driving_3_lines.png. The amplitude of these oscilations seem to be damped ~15 dB. There seems to be an additional peak around 450 Hz that is more broadband that should be investigated.

Jeff had an idea to play some white noise and try the same thing. Result can be found in white_noise_injection.png. Here we can see the resonance is completely damped out. It should be noted that the amplitude of the white noise the phone app I was using was quite small, so this is probably just damping out the oscillations due to the ambient noise of the room. Daniel found a louder app, will repeat this with bigger white noise. We should buy a speaker and microphone for future tests. Data for these can be found in the zip.

I find this result quite strange as we are only damping one mirror. I am going to repeat this process doing the same thing to M4. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 16:44, Friday 19 January 2024 (11428)

Touching M4 at all misaligns the cavity enough to not longer see any transmission peaks at all. It seems like that mirror in the flexture is much less rigidly held. Cannot perform this test with M4

Equipment General
Ian MacMillan - posted 12:30, Friday 19 January 2024 (11426)
Acquiring Solvents

[Ian, Daniel]

We went to the Caltech VWR Stockroom to get solvents. We got

To get them we went to the stockroom linked above, filled out the PTA number, and loaded them onto the cart.

We still need to buy funnels to fill our smaller bottles. we need to buy appropriate funnels such that they are safe for these chemicals. We have a bottle for Methanol but we decided to hold off on purchasing that until later as we don't really need it yet. 

We need to move them to a chemical cabinet. Right now they are stored in a fume hood that is not turned on I want to temporarily store them in one of the chemical cabinets (see image) in one of the Adhikari labs. I want this to happen today so I will ask for space ASAP.

Controls General (GQuEST)
Chris Stoughton - posted 08:36, Friday 19 January 2024 (11425)
DMA on RedPitaya

Another minor step towards pynq on the RedPitaya control system.  Here is the result of uploading an arbitrary wave form, playing it, and reading it back.  This is the demo described here:  https://github.com/dspsandbox/FPGA-Notes-for-Scientists/wiki/DMA-transfer

 

One vhdl bug had to be fixed, putting a 0 value in quotes to help the compiler.  Vivado compilers are more careful now about types.

Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:26, Thursday 18 January 2024 - last comment - 07:07, Friday 19 January 2024(11423)
mystery resonance investigation

[Daniel, Mai ,Torrey]

The leading theory on the ~520 Hz resonance seen in the spectrum of the cavity error signal is some part of the flexture mount mirrors resonating. To test this theory we pushed on the mirror with viton to see if the resonance is damped/moved at all. Our initial test has a viton O ring pushing on M2 of the cavity only. We did this with and without pushing on the mirror, results seen here image.png. We are skeptical of these results and tomorrow will:

a) repeat this test several times. It is time consuming as the force on the cavity mirrors misaligns the cavity.

b) instead of relying on the ambient noise of the room to ring up, use phones to drive at say 520 +/- 20 Hz at constant locations between tests.

Images attached to this report
Comments related to this report
Chris Stoughton - 07:07, Friday 19 January 2024 (11424)

Can you rig up something to do a swept sine of a speaker placed on the table?  This will probably be enought to drive the resonance.

When you "push with an o-ring" I think that will absorb energy from vibrational modes.  Is that correct?

 

Is it possible to have the entire cavity sit on viton?  A quick way to do this is to have a sheet of viton between the cavity and the table.  

We shipped out some of the "platforms" that were inside the Holometer vacuum systems.  These have recesses for viton balls so the platform is stable. I do not recall the cutoff frequencies of that system, but it eliminated our acoustic problems.

GQuEST Output Filter Cavities
Torrey Cullen - posted 15:24, Thursday 18 January 2024 (11422)
further cavity science

[Torrey, Lee, Daniel]

After some investigation we have found a mystery resonance around ~520 Hz. IMG_0015.png is the set up used to take a spectrum of the error signal and the control signal. As seen from mysteryresonance.png (blue is error, orange control) you can see a resonance peak at ~520. This exibits resonance type behavior, i.e. drive a sound at 520 Hz and measure the level of excitation and compare it to frequencies around this, say +/- 20 hz it is much reduced. 

Additionally we found some source of phase delay from our last post by taking a transfer function accros the piezo controller, as seen in IMG_0020.png. Lee mentioned it would be useful to create a filter that inverts this amplitude loss.

Images attached to this report
Non-image files attached to this report
General (GQuEST)
Chris Stoughton - posted 08:13, Thursday 18 January 2024 (11420)
RedPitaya analog_echo

The same sd card in a different RedPitaya and power supply works.  

 

Yellow is the input signal to the RedPitays.  Blue is the output, with offset register set to 100 mV.  The delay seems to be 100 ns.  The gain seems to be a bit more than 1.

 

So, we can read and write analog signals and control registers with pynq.

 

 

Images attached to this report
GQuEST General (GQuEST)
Daniel Grass - posted 17:17, Wednesday 17 January 2024 - last comment - 14:19, Thursday 18 January 2024(11418)
Power Distribution Sled Power Drift Test

[Torrey, Daniel]

We aligned light from the misc category into a fiber on the power distribution sled. Instead of the 5 axis mount, we used a simple 2 axis mirror mount. We also tuned up the alignment into the other 3 fibers to maximize the power. We measured the power going into the fiber and out of the fiber to see if and how it drifts over time. Potentially, the 5 axis mounts will drift more than the 2 axis mount.

Comments related to this report
Chris Stoughton - 07:09, Thursday 18 January 2024 (11419)

Can you say what the "misc category" means?  Thank you.

Torrey Cullen - 14:19, Thursday 18 January 2024 (11421)

Hey Chris. The power distribution sled was designed to have 4 paths, 3 of the paths being ear marked for the output filter cavity, laser filter cavity, and SHG sleds. This last path is just for miscellaneous needs. Its going into a fiber that can be piped to any point in the room where we might need an extra laser source.

Controls General
Chris Stoughton - posted 12:35, Wednesday 17 January 2024 (11417)
RedPItaya -- analog feedback

I built the example analog feedback from https://github.com/dspsandbox/Pynq-Redpitaya-125

 

On one RedPitaya board the firmare loads, but setting the offset value via pynq causes the board to freeze up.  Identical firmaware and test on a second board succeeds.  Now will diagnose the "bad" board.  First suspect is the power supply.

Controls General (GQuEST)
Lee McCuller - posted 09:11, Tuesday 16 January 2024 - last comment - 23:01, Wednesday 20 November 2024(11413)
Filter cavity control

[Lee, Torrey, Daniel]

On Friday, we adjusted the loop shaping on the filter cavity to try and make the lock better. With a 1/f loop, it will lock, but goes between flickering out of lock, to oscillating as the gain is increased.

The new shaping uses the "slow" output of the "laser lock box." It shapes the fast output to have a 10Hz - 200Hz boost (pole - zero), giving 20x gain at low frequencies. The overall 1/f is then put at 1Hz in the slow path. We are able to get a UGF of 1kHz, with 30-40deg remaining. We lose 90deg by 2kHz from either the actuator or the Moku, so 1kHz is around the max UGF we can currently get.

Even with that loop, we get dropouts while talking loudly or while the fans are on. Problematic. That mean the loop can supress the peak length noise below 1.5um / 2  / (2 * finesse). Our current input coupler is 1%, so our finesse is around 2pi / 0.01 = 600 (it might be as low as 300, depending on the losses in the cavity). So our peaks/RMS are around 0.5nm - 1nm of noise.

 

+1 to the piezo transfer teams. both 11380 and 11373 have the transfer function data along with the plots. It looks like all of the phase we are losing in the cavity lock is from the piezo (though could still be from Moku). We'll need to fit and at least slightly invert this actuator transfer function to make a loop above 1kHz UGF. The fit can help tell us how much of the phase is from the piezo roll-off.

We are going to need it - the current loop & cavity  has marginally too much noise, with an RMS around the linewidth. That is going to be a problem once we increase the finesse of the cavity. So we'll need to reduce the noise and/or improve the loop a respectable amount.

our voices are driving 100-300Hz, so we need gain there. If we can get the UGF up as high as 3kHz, then we should also be able to move the boosts up to 600Hz or so. That should give 9x more loop gain and cover much of the need.

options:

frequency lock - good for acquiring and transitioning to aggressive lock, but not an option for operating the experiment

optimal control - we can determine how good we can do with what we have

fitting/inverting actuator - probably needed to get UGF higher. May vary with time in bad ways

sound attenuation or active noise cancellation - may be possible and should look into engineering this.

 

We should determine if the sound driving is from pushing the mirrors or from the index fluctuation as the pressure modulates. 11290 has calculations about the pressure pushing on a 1kg mass. My worry is that the small alignment adjust set-screws in the MKS flexture mounts are fairly compliant. Could sound drive them 1nm?
My guess is that the light going through 2m of air is the culprit.

The sound pressure of 40dB should be around 2000uP. Air pressure is 101kPa. The index of air is 1.000293

2m / (1.000293**2) * 2000e-6Pa / 101kPa * (1.000293 - 1) = 1.2e-11m

So that is a little small, but if our sound is at 80db, which would be quite loud, then it would be enough to cause this.

For the mirror motion, the mirror mounts are about 1.5in square. On the pressure level that is

(.0254m * 1.5)**2 * 2000e-6Pa = 3e-6N

of the mirror+panel is about 35g (very rough estimate). For a free mass motion that is

1 / ((2 * np.pi * 100Hz)**2 * 0.03534291735288516kg) * 3e-6N = 2e-10m which is getting in the ballpark with 60db of noise rather than 40db. The mirrors aren't free though.

and now I'm curious how that is against the spring constant of the tiny screw.

Comments related to this report
Torrey Cullen - 10:49, Tuesday 16 January 2024 (11414)

Supplement to Lee's post:

1) IMG_0005.png is the scheme used to achieve lock in multi-instrument mode.

2) IMG_0004.png is the closed loop gain while the cavity is locked. Note the UGF of ~1kHz as Lee mentioned.

3) IMG_0009.png is the slow controller loop.

4) IMG_0008.pngis a spectrum for f<10kHz. Note the resonance at ~3kHz. This may be the culprit and will be investigated.

5) IMG_0007.png is the fast controller loop just being used as a low frequency boost.

6) IMG_0010(1).png is the overall structure of the laser lock box as a reference for people that build this system in the future.

Images attached to this comment
Torrey Cullen - 11:28, Tuesday 16 January 2024 (11415)

Quick test to see if the air theory is likely. IMG_0011.png is cavity locked, fans off, no noise emitting as a control. IMG_0014.png is cavity locked, emitting a 1kHz sound from my phone about 1.5 meters away from the cavity with the lid off (reference trace) and lid on (active measurement). Noise is definitely muted. Air theory seems plausible.

Images attached to this comment
Daniel Grass - 14:10, Tuesday 16 January 2024 (11416)

Flexure Mount spring constant: E*A/L = 200 GPA * pi * 0.035 in ^2 / 0.1 in = 2*10^8 N/m

Resonant frequency: 75 kHz

RMS displacement: 2 * 10^-14 m for 40 dB of sound

Daniel Grass - 23:01, Wednesday 20 November 2024 (12007)

The small piece of metal that flexes on this mount is 0.020" thick and 1.5" long according to SolidWorks (this seems about right experimentally). This is a cross section 10x of the screw, so this also doesn't explain the acoustic pick-up.

Equipment General
Daniel Grass - posted 11:45, Wednesday 10 January 2024 (11411)
Newport 10D20DM.8 (1550 nm HR mirror) reflectivity of 775 nm light

[Torrey, Daniel]

We measured the reflectivity of 775 nm light on the Newport 10D20DM.8 (1550 nm HR mirror). At a 45° AOI, there is roughly a 40-50% reflection. +/- around 5° away from 45°, the reflection drops off sharply. Thus, these mirrors cannot be used to make a bowtie cavity for 775 nm light.

Low AOI: 3% reflection, 70% "transmission" but definitely a lot of this is scattered light

45°: 40-50% reflection depending on polarization, ~40% "transmission" but definitely a lot of this is scattered light

SNSPD General
Chris Stoughton - posted 07:02, Tuesday 09 January 2024 (11410)
SNSPD Dewar taking shape

The frame is assembled.  Custom-machined parts for the walls and inside shields are on order.

 

 

Images attached to this report
Displaying reports 861-880 of 1079.Go to page Start 40 41 42 43 44 45 46 47 48 End